QUESTION 3

If the alternative hypothesis has a mathematical sign of not equal (≠) and the experimental Z value is +3.0 then the p – value is 0.0027.

True

False

QUESTION 4

When the mean is 52, the sample size is 25, and the standard deviation is 2, the probability of committing Type II error when P( 50 <    < 54) is one (1).

True

False
Answer #1

3)

p value for above test statistic and two tailed test =2*P(Z>3) =2*0.00135 =0.0027

true

4)

sample size       =n= 25
std error=σ=σ/√n= 0.40
probability =P(50<X<54)=P((50-52)/0.4)<Z<(54-52)/0.4)=P(-5<Z<5)=1-0=1.0000

true